OpenStudy (anonymous):

factor completely 25-49v^2

OpenStudy (blurbendy):

it would good start by factoring out the negative sign - ( ) ( ) this will make it easier to factor. It's safe to say the answer will involve a 7 and a 5 (because 7 * 7 = 49 and 5 * 5 = 25 -(7v - 5) ( ) Seems promising so far. Can you fill in what goes between the other parentheses?